The perimeter of rectangle is 86 inches. The width of the rectangle is 5 less than two times its length. What is the area of the rectangle?
Answer by TimothyLamb(4379)

---
2x + 2y = 86
y = 2x - 5
---
put the system of linear equations into standard form:
---
2x + 2y = 86
2x - y = 5
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 16
y = 27
---
area = xy
area = 16*27
area = 432 sq.in
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php